# Ask in Any Modality: A Comprehensive Survey on Multimodal Retrieval-Augmented Generation

### v2

@@ Appendix C Metrics
FID (Fréchet inception distance): FID is a metric used to assess the quality of images created by a generative model. The formula for FID is:
where $\mu_{r}$ and $\Sigma_{r}$ are the mean and covariance of real images’ feature representations, respectively. $\mu_{g}$ and $\Sigma_{g}$ are the mean and covariance of generated images’ feature representations, respectively. To extract features, InceptionV3 is typically used.
ROUGE-N (N-gram Recall): The ROUGE metric is commonly used to evaluate text summarization and generation. ROUGE-N measures the overlap of N-grams between the generated and reference text. The formula for ROUGE-N is:
@@
BLEU: BLEU is another metric used for assessing text generation. The formula for calculating BLEU is:
@@
BERTScore: BERTScore is a metric for evaluating the quality of text generation, based on the similarity between the contextual embeddings of words in the candidate and reference texts. The formula for calculating BERTScore is:
$\text{BERTScore}(c,r)=\frac{1}{|c|}\sum_{i=1}^{|c|}\max_{j=1}^{|r|}\cos(\mathbf{e}_{i},\mathbf{e}_{j})$ (9)
$c$ is the candidate sentence, and $r$ is the reference sentence. $e_{i}$ and $e_{j}$ are the embeddings (e.g., from BERT) for words $c_{i}$ and $r_{j}$ in the candidate and reference sentences, respectively.
CLIPScore: CLIPScore is a metric that evaluates the similarity between the text and an image by using the CLIP model. The formula for calculating CLIPScore is:
$\text{CLIPScore}=\frac{\cos(\mathbf{t},\mathbf{i})}{\|\mathbf{t}\|\cdot\|\mathbf{i}\|}$ (10)
where t and i are text and image embedding respectively.
Mean Reciprocal Rank (MRR): MRR is a metric used to evaluate the performance of systems that return a list of results, such as search engines or recommendation systems. MRR measures the rank position of the first relevant result in the returned list. The formula for calculating MRR is:
$\text{MRR}=\frac{1}{Q}\sum_{q=1}^{Q}\frac{1}{\text{rank}_{q}}$ (11)
Appendix D Loss Function
@@
where $x_{a}^{i}$ is the anchor sample. $x_{p}^{i}$ and $x_{n}^{i}$ are the positive and negative samples respectively. $f(x)$ is the neural network.

### v3

@@ Appendix C Evaluation and Metrics
Recall@K, which examines relevant items in top K results, is preferred over standard recall. Mean Reciprocal Rank (MRR) serves as another key metric for evaluation, which is utilized by Adjali et al. (2024); Nguyen et al. (2024). MRR measures the rank position of the first relevant result in the returned list. The formula for calculating MRR is:
$\text{MRR}=\frac{1}{Q}\sum_{q=1}^{Q}\frac{1}{\text{rank}_{q}}$ (2)
Modality Evaluation
Modality-based evaluations primarily focus on text and image, assessing their alignment, text fluency, and image caption quality. For text evaluation, metrics include Exact Match (EM), BLEU Papineni et al. (2002), ROUGE Lin (2004), and METEOR Banerjee and Lavie (2005). The ROUGE metric is commonly used to evaluate text summarization and generation. ROUGE-N measures the overlap of N-grams between the generated and reference text. The formula for ROUGE-N is:
@@
BLEU is another metric used for assessing text generation. The formula for calculating BLEU is:
@@
MultiRAGen Shohan et al. (2024) uses Multilingual ROUGE for multilingual settings.
For image captioning, CIDEr (Consensus-Based Image Description Evaluation) Vedantam et al. (2015) measures caption quality using TF-IDF and cosine similarity Yasunaga et al. (2023); Zhao et al. (2024); Luo et al. (2024a); Yuan et al. (2024); Sharifymoghaddam et al. (2024); Hu et al. (2023); Rao et al. (2024); Xu et al. (2024a); Kim et al. (2024); Zhang et al. (2024c), while SPICE (Semantic Propositional Image Caption Evaluation) Anderson et al. (2016) focuses on semantics. SPIDEr Liu et al. (2017), used in Zhang et al. (2024c), combines both metrics.
For semantic alignment, BERTScore Zhang et al. (2020) compares BERT embeddings Sun et al. (2024b); Shohan et al. (2024), and evaluates fluency Chen et al. (2022a); Zhi Lim et al. (2024); Ma et al. (2024d).
CLIP Score Hessel et al. (2021), used in Sharifymoghaddam et al. (2024); Zhang et al. (2024c), measures image-text similarity using CLIP Radford et al. (2021). The formula for calculating CLIPScore is:
$\text{CLIPScore}=\frac{\mathbf{t}.\mathbf{i}}{\|\mathbf{t}\|\|\mathbf{i}\|}$ (7)
where t and i are text and image embedding, respectively.
For image quality, FID (Fréchet Inception Distance) Heusel et al. (2017) compares feature distributions Yasunaga et al. (2023); Zhao et al. (2024); Sharifymoghaddam et al. (2024); Zhang et al. (2024c), while KID (Kernel Inception Distance) Bińkowski et al. (2018) provides an unbiased alternative. The formula for FID is:
where $\mu_{r}$ and $\Sigma_{r}$ are the mean and covariance of real images’ feature representations, respectively. $\mu_{g}$ and $\Sigma_{g}$ are the mean and covariance of generated images’ feature representations, respectively. To extract features, InceptionV3 Szegedy et al. (2016) is typically used.
Inception Score (IS) evaluates image diversity and quality through classification probabilities Zhi Lim et al. (2024). For audio evaluation, Zhang et al. (2024c) use human annotators to assess sound quality (OVL) and text relevance (REL), while also employing Fréchet Audio Distance (FAD) Kilgour et al. (2019), an audio-specific variant of FID.
Appendix D Robustness Advancements and Loss Functions
D.1 Robustness and Noise Management
Multimodal training faces challenges such as noise and modality-specific biases Buettner and Kovashka (2024). Managing noisy retrieval inputs is critical for maintaining model performance. MORE Cui et al. (2024) injects irrelevant results during training to enhance focus on relevant inputs. AlzheimerRAG Lahiri and Hu (2024) uses progressive knowledge distillation to reduce noise while maintaining multimodal alignment. RAGTrans Cheng et al. (2024) leverages hypergraph-based knowledge aggregation to refine multimodal representations, ensuring more effective propagation of relevant information. RA-BLIP Ding et al. (2024b) introduces the Adaptive Selection Knowledge Generation (ASKG) strategy, which leverages the implicit capabilities of LLMs to filter relevant knowledge for generation through a denoising-enhanced loss term, eliminating the need for fine-tuning. This approach achieves strong performance compared to baselines while significantly reducing computational overhead by minimizing trainable parameters. RagVL Chen et al. (2024e) improves robustness through noise-injected training by adding hard negative samples at the data level and applying Gaussian noise with loss reweighting at the token level, enhancing the model’s resilience to multimodal noise. Finally, RA-CM3 Yasunaga et al. (2023) enhances generalization using Query Dropout, which randomly removes query tokens during retrieval, serving as a regularization method that improves generator performance.
D.2 Loss Function
@@
where $x_{a}^{i}$ is the anchor sample. $x_{p}^{i}$ and $x_{n}^{i}$ are the positive and negative samples, respectively. $f(x)$ is the neural network.
Appendix E Applications and Relevant Tasks
Multimodal RAG extends traditional RAG beyond unimodal settings to cross-modal tasks. In content generation, it enhances image captioning Zhi Lim et al. (2024); Hu et al. (2023); Rao et al. (2024) and text-to-image synthesis Yasunaga et al. (2023); Chen et al. (2022b) by retrieving relevant contextual information. It also improves coherence in visual storytelling and factual alignment in multimodal summarization Tonmoy et al. (2024). In knowledge-intensive applications, multimodal RAG supports open-domain QA Chen et al. (2024e); Ding et al. (2024b); Yuan et al. (2023), video-based QA Luo et al. (2024b), fact verification Khaliq et al. (2024), and zero-shot image–text retrieval Yang et al. (2024b), grounding responses in retrieved knowledge and thereby mitigating hallucinations.
Additionally, the incorporation of chain-of-thought reasoning Zhai (2024); Khaliq et al. (2024) further enhances complex problem-solving and inference. Finally, their integration into AI assistants such as Gemini Team et al. (2024) enables natural language-driven visual search, document understanding, and multimodal reasoning.
Multimodal RAGs are increasingly applied across diverse domains, including healthcare, software engineering, and creative industries (e.g., fashion and design automation). The taxonomy of application domains can be seen in  Figure 3. The following sections explore domain-specific adaptations of these techniques in greater depth.
Healthcare and Medicine
Multimodal RAG enhances clinical decision-making through integrated analysis of medical imaging, electronic health records, and biomedical literature. Systems like MMED-RAG Xia et al. (2024a) address diagnostic uncertainty in medical visual question answering by aligning radiology images with contextual patient data. In automated report generation, RULE Xia et al. (2024b) mitigates hallucinations through dynamic retrieval of clinically similar cases. Similarly, RA-RRG Choi et al. (2025) first leverages an LLM to extract key textual phrases from a report corpus, then employs a multimodal retriever to link the visual features to these relevant phrases. The coherent report is generated after being retrieved by another LLM without fine-tuning, thereby reducing hallucinations. FactMM-RAG Sun et al. (2024b) further automates radiology report drafting by retrieving biomarker correlations from medical ontologies, exemplifying the paradigm’s capacity to operationalize expert knowledge at scale. AsthmaBot Bahaj and Ghogho (2024) introduces a multimodal RAG-based approach for supporting asthma patients across multiple languages, enabling structured, language-specific semantic searches. Predictive frameworks such as Realm Zhu et al. (2024c) demonstrate robust risk assessment by fusing heterogeneous patient data streams, while Su et al. (2024a) advance privacy-preserving architectures for federated clinical data integration.
Software Engineering
Fashion and E-Commerce
Cross-modal alignment drives advancements in product discovery and design automation. UniFashion Zhao et al. (2024) enables style-aware retrieval by jointly embedding garment images and textual descriptors, while Dang (2024) reduces search friction through multimodal query expansion. For fashion image editing, Fashion-RAG Sanguigni et al. (2025) employs a retrieval-augmented approach, retrieving garments that match textual descriptions and integrating their attributes into image generation via textual inversion techniques within diffusion models, ensuring personalized and contextually relevant outputs. LLM4DESIGN Chen et al. (2024d) demonstrates architectural design automation by retrieving compliance constraints and environmental impact assessments, underscoring RAG’s adaptability to creative domains.
Entertainment and Social Computing
Emerging Applications
Appendix F Additional Future Directions
High computational costs in video frame sampling and memory bottlenecks in processing multi-page documents with images remain key challenges in long-context processing. Fixed extraction rates struggle to capture relevant frames, requiring adaptive selection based on content complexity and movement Kandhare and Gisselbrecht (2024). Additionally, retrieval speed-accuracy trade-offs in edge deployments and redundant computations in cross-modal fusion layers emphasize the need for efficient, scalable architectures. Personalization mechanisms, like user-specific retrieval (e.g., adapting to medical history), remain underexplored. As these systems evolve, ensuring privacy and preventing sensitive data leakage in multimodal outputs is critical. Lastly, the lack of datasets with complex reasoning tasks and multimodal adversarial examples limits robust evaluation.